# What is the weakest idempotent Maltsev condition that implies that abelian tolerances generate abelian congruences?

### v1

@@ 1. Introduction
    Whenever ${\mathbf{A}}\in\mathcal{V}$ and $\alpha\in\operatorname{Con}({\mathbf{A}})$, there is no pentagon labeled as in Figure 1 with $[\alpha,\alpha]=0$. (No “spanning pentagon” in $I[0,\alpha]$.)
@@
    Whenever ${\mathbf{A}}\in\mathcal{V}$ and $\alpha\in\operatorname{Con}({\mathbf{A}})$ is abelian, there is no pentagon labeled as in Figure 1 where $[\alpha,\alpha]=0$ and ${\mathbf{C}}(\theta,\alpha;\delta)$.
    $\beta$$\delta$$\theta$$\alpha$00
    Figure 1. Forbidden sublattice if both $[\alpha,\alpha]=0$ and ${\mathbf{C}}(\theta,\alpha;\delta)$ hold.
@@ 2. What is the error?
and we could indeed conclude from Theorem 2.19 (8) that ${\mathbf{C}}(R_{i},\Delta;0)$ holds, and the proof could be concluded as we did in [3]. Unfortunately, in the proof of Theorem 3.24 of [3], $R_{i}$ is only a tolerance and not necessarily a congruence (it is a reflexive, transitive, compatible, binary relation, but it need not be a transitive relation). This means that Theorem 2.19 (8) does not apply to the situation. Theorem 2.19 (7) is more appropriate for tolerances. Using it, one may correctly derive from Claim 3.25 (i.e., from the claim that $\Delta\cap R_{i}=0$ holds) that ${\mathbf{C}}(\Delta,R_{i};0)$ holds. But this is different than ${\mathbf{C}}(R_{i},\Delta;0)$, which is what we needed for the rest of the proof of Theorem 3.24. If we had proved a slightly stronger statement in Claim 3.25, namely that $\Delta\cap(R_{i}\circ R_{i})=0$, then we could correctly derive from Theorem 2.19 (7) that ${\mathbf{C}}(R_{i},\Delta;0)$ holds. But, as it happens, there is no way to correct the proof of Theorem 3.24, since the statement of the theorem is false. It will be corrected and strengthened in Theorem 3.1 of the next section.
@@ 3. The Main Correction
In more detail, a Taylor term for a variety $\mathcal{V}$ is a $\mathcal{V}$-term $T(x_{1},\ldots,x_{n})$ such that $\mathcal{V}$ satisfies the identity $T(x,\ldots,x)\approx x$ and, for each $i$ between $1$ and $n$, $\mathcal{V}$ satisfies some identity of the form $T({\mathbf{w}})\approx T({\mathbf{z}})$ with $w_{i}\neq z_{i}$. A weak difference term for $\mathcal{V}$ is a $\mathcal{V}$-term $w(x,y,z)$ such that for any ${\mathbf{B}}\in{\mathcal{V}}$, $w^{{\mathbf{B}}}(a,a,b)=b=w^{{\mathbf{B}}}(b,a,a)$ holds whenever the pair $(a,b)$ is contained in an abelian congruence. It is known that the class of varieties with a weak difference term is definable by a nontrivial idempotent Maltsev condition (see [4, Theorem 4.8]). It is known that the class of varieties with a Taylor term is definable by the weakest nontrivial idempotent Maltsev condition. (This is derivable from Corollaries 5.2 and 5.3 of [6].) In particular, this proves that ${\mathscr{W}}\Rightarrow{\mathscr{T}}$. It is known that ${\mathscr{T}}\not\Rightarrow{\mathscr{W}}$ for arbitrary varieties (see [2, Example 4.4]). It is also known that ${\mathscr{W}}\Leftrightarrow{\mathscr{T}}$ for locally finite varieties (see [1, Theorem 9.6]). These facts, (i) ${\mathscr{W}}\Rightarrow{\mathscr{T}}$, (ii) ${\mathscr{T}}\not\Rightarrow{\mathscr{W}}$, (iii) ${\mathscr{W}}\Leftrightarrow{\mathscr{T}}$ for locally finite varieties, are the data supporting our statement that “$\mathscr{W}$ is slightly (but properly) stronger than $\mathscr{T}$”.
@@
Now assume that Item (2) holds. Choose any ${\mathbf{A}}\in{\mathcal{V}}$ and any abelian tolerance $T$ of ${\mathbf{A}}$. Our goal is to prove that $T$ is transitive, so choose elements satisfying $a\;T\;b\;T\;c$. Since $(a,b),(b,c)\in T$ and $T$ is a reflexive subalgebra of ${\mathbf{A}}^{2}$ we have $w^{T}((a,b),(b,b),(b,c))\in T$. Since $w(x,y,z)$ is a weak difference term for tolerances and $T$ is an abelian tolerance containing $(a,b)$ and $(b,c)$, we have $w^{T}((a,b),(b,b),(b,c))=(w^{{\mathbf{A}}}(a,b,b),w^{{\mathbf{A}}}(b,b,c))=(a,c)$, proving that $(a,c)\in T$. This shows that Item (3) holds.
The implication (3$\Rightarrow$(4) holds since the congruence generated by a congruence is itself.
To prove the final implication (4$\Rightarrow$(1) we explain instead why $\neg(1)\Rightarrow\neg(4)$ holds. Assume that Item (1) fails. According to Theorem 4.5 of [2], which is stated in the Introduction of this article, some ${\mathbf{A}}\in\mathcal{V}$ must have a pentagon in its congruence lattice labeled as in Figure 1, and for this pentagon we may assume that $[\alpha,\alpha]=0$ and ${\mathbf{C}}(\theta,\alpha;\delta)$. Let $T=(\delta\circ\beta\circ\beta)/\delta$. We will complete the proof that (4) fails by arguing that $T$ is an abelian tolerance on ${\mathbf{A}}/\delta$ which does not generate an abelian congruence on ${\mathbf{A}}/\delta$.
For the congruences depicted in Figure 1, we have $[\alpha,\alpha]=0$ and $\beta\leq\alpha$, so $[\beta,\beta]=0$ by monotonicity of the commutator. In terms of the centralizer relation, this means ${\mathbf{C}}(\beta,\beta;0)$ holds. Now ${\mathbf{C}}(\beta,\beta;0)$ is equivalent to ${\mathbf{C}}(\beta,\beta;\delta)$ by Theorem 2.19 (4) above, so ${\mathbf{C}}(\beta,\beta;\delta)$ must also hold. We also have ${\mathbf{C}}(\delta,\beta;\delta)$ by Theorem 2.19 (8). The statements ${\mathbf{C}}(\beta,\beta;\delta)$ and ${\mathbf{C}}(\delta,\beta;\delta)$ may be combined to ${\mathbf{C}}(\beta\vee\delta,\beta;\delta)$ by Theorem 2.19 (5). By Theorem 2.19 (3) we derive that ${\mathbf{C}}(\beta\vee\delta,\delta\circ\beta\circ\delta;\delta)$ holds. If $S$ is the tolerance $\delta\circ\beta\circ\delta$ on ${\mathbf{A}}$, then since $S\subseteq\beta\vee\delta$ we derive from ${\mathbf{C}}(\beta\vee\delta,\delta\circ\beta\circ\delta;\delta)$ and Theorem 2.19 (1) that ${\mathbf{C}}(S,S;\delta)$ holds. Now, by Theorem 2.19 (10), we derive that ${\mathbf{C}}(S/\delta,S/\delta;0)$ holds. But $T=S/\delta$, so ${\mathbf{C}}(T,T;0)$ holds. We have proved that $T$ is an abelian tolerance of ${\mathbf{A}}/\delta$.
@@
Theorem 3.24 of [3] is correct if the hypothesis that $\mathcal{V}$ has a Taylor term (or “$\mathcal{V}$ satisfies a nontrivial idempotent Maltsev condition”) is strengthened to the hypothesis that $\mathcal{V}$ has a weak difference term.
@@ 4. Consequences
    ([3, Theorem 3.27]) This theorem examines the following property of a variety $\mathcal{V}$: given two perspective congruence intervals $I[\alpha,\alpha\wedge\beta]$ and $I[\alpha\vee\beta,\beta]$ of some ${\mathbf{A}}\in\mathcal{V}$, one interval is abelian if and only the other is abelian. Call this property $\mathscr{B}$. [3, Theorem 3.27] asserts that ${\mathscr{T}}\Rightarrow{\mathscr{B}}$. This implication is false. Using Corollary 3.2, we can correct Theorem 3.27 by strengthening the hypothesis $\mathscr{T}$ to the hypothesis $\mathscr{W}$. The corrected version then expresses that ${\mathscr{W}}\Rightarrow{\mathscr{B}}$. In fact, it is not hard to see that ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{B}})$.
@@
    • All claims of all items in Theorem 4.16 are true when the hypothesis “${\mathcal{V}}$ has a Taylor term” is strengthened to “${\mathcal{V}}$ has a weak difference term”. (However, the two claims of 4.16 (3) were already known to be true in the presence of a weak difference term, cf. [5, Corollary 5.8].)
    We give more detail about the second bullet point. Using the labeling in Figure 1, it is stated in [3, Theorem 4.16 (2)] that a pentagon of congruences where ${\mathbf{C}}(\beta,\theta;\delta)$ holds cannot occur as a sublattice in a variety with a Taylor term. The proof given for this is correct. But then it is stated that a pentagon of congruences where ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ holds cannot occur as a sublattice. Call this property ${\mathscr{C}}$. The proof given that ${\mathscr{T}}\Rightarrow{\mathscr{C}}$ depends on [3, Theorem 3.27], which (we have seen) is false without a weak difference term. In fact, it can be shown by arguments like those above that ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{C}})$. To summarize, the part of [3, Theorem 4.16 (2)] that refers to ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ is false as stated, but becomes true when the hypothesis “${\mathcal{V}}$ has a Taylor term” is strengthened to “${\mathcal{V}}$ has a weak difference term”.
@@
    ([3, Theorem 8.1]) This theorem refers to [3, Theorem 3.24] in its proof of (4$\Rightarrow$(5). The reference can be eliminated, as we now explain. Item (4) is the assertion that $\mathcal{V}$ contains no algebra with a nontrivial abelian congruence. Item (5) is the assertion that $\mathcal{V}$ contains no algebra with a nontrivial abelian tolerance. What we need to establish, therefore, is that any variety which omits nontrivial abelian congruences also omits nontrivial abelian tolerances. We can use Theorem 3.1 above to do this. Assume that Item (4) of [3, Theorem 8.1] holds. Since $\mathcal{V}$ contains no algebra with a nontrivial abelian congruence, the ternary projection $w(x,y,z):=x$ is a weak difference term for $\mathcal{V}$. Any variety with a weak difference term has a Taylor term, so Theorem 3.1 applies to $\mathcal{V}$. In particular, from the implication (1$\Rightarrow$(3) of Theorem 3.1 we see that an abelian tolerance on an algebra in $\mathcal{V}$ is a congruence. Since abelian congruences in $\mathcal{V}$ are trivial, abelian tolerances in $\mathcal{V}$ are also trivial, so Item (5) of [3, Theorem 8.1] holds. This shows that [3, Theorem 8.1] is true as stated.
@@
    ([3, Theorem 3.27]) is false as stated. Theorem 3.27 is true when the hypothesis “${\mathcal{V}}$ has a Taylor term” is strengthened to “${\mathcal{V}}$ has a weak difference term”.
@@
    • All claims of all items in Theorem 4.16 are true when the hypothesis “${\mathcal{V}}$ has a Taylor term” is strengthened to “${\mathcal{V}}$ has a weak difference term”.

### v2

@@ 1. Introduction
    Whenever ${\mathbf{A}}\in\mathcal{V}$ and $\alpha\in\operatorname{Con}({\mathbf{A}})$, there is no pentagon labeled as in Figure 1 with $\alpha,\alpha]=0$. (No “spanning pentagon” in $I[0,\alpha]$.)
@@
    Whenever ${\mathbf{A}}\in\mathcal{V}$ and $\alpha\in\operatorname{Con}({\mathbf{A}})$ is abelian, there is no pentagon labeled as in Figure 1 where $\alpha,\alpha]=0$ and ${\mathbf{C}}(\theta,\alpha;\delta)$.
    $\beta$$\delta$$\theta$$\alpha$0
    Figure 1. Forbidden sublattice if both $\alpha,\alpha]=0$ and ${\mathbf{C}}(\theta,\alpha;\delta)$ hold.
@@ 2. What is the error?
and we could indeed conclude from Theorem 2.19 (8) that ${\mathbf{C}}(R_{i},\Delta;0)$ holds, and the proof could be concluded as we did in [3]. Unfortunately, in the proof of Theorem 3.24 of [3], $R_{i}$ is only a tolerance and not necessarily a congruence (it is a reflexive, symmetric, compatible, binary relation, but it need not be a transitive relation). This means that Theorem 2.19 (8) does not apply to the situation. Theorem 2.19 (7) is more appropriate for tolerances. Using it, one may correctly derive from Claim 3.25 (i.e., from the claim that $\Delta\cap R_{i}=0$ holds) that ${\mathbf{C}}(\Delta,R_{i};0)$ holds. But this is different than ${\mathbf{C}}(R_{i},\Delta;0)$, which is what we needed for the rest of the proof of Theorem 3.24. If we had proved a slightly stronger statement in Claim 3.25, namely that $\Delta\cap(R_{i}\circ R_{i})=0$, then we could correctly derive from Theorem 2.19 (7) that ${\mathbf{C}}(R_{i},\Delta;0)$ holds. But, as it happens, there is no way to correct the proof of Theorem 3.24, since the statement of the theorem is false, as we will see. It will be corrected and strengthened in Theorem 3.1 of the next section.
@@ 3. The Main Correction
In more detail, a Taylor term for a variety $\mathcal{V}$ is a $\mathcal{V}$-term $T(x_{1},\ldots,x_{n})$ such that $\mathcal{V}$ satisfies the identity $T(x,\ldots,x)\approx x$ and, for each $i$ between $1$ and $n$, $\mathcal{V}$ satisfies some identity of the form $T({\mathbf{w}})\approx T({\mathbf{z}})$ with $w_{i}\neq z_{i}$. A weak difference term for $\mathcal{V}$ is a $\mathcal{V}$-term $w(x,y,z)$ such that for any ${\mathbf{B}}\in{\mathcal{V}}$, $w^{{\mathbf{B}}}(a,a,b)=b=w^{{\mathbf{B}}}(b,a,a)$ holds whenever the pair $(a,b)$ is contained in an abelian congruence. It is known that the class of varieties with a weak difference term is definable by a nontrivial idempotent Maltsev condition (see [4, Theorem 4.8]). It is known that the class of varieties with a Taylor term is definable by the weakest nontrivial idempotent Maltsev condition. (This is derivable from Corollaries 5.2 and 5.3 of [6].) In particular, this proves that ${\mathscr{W}}\Rightarrow{\mathscr{T}}$. It is known that ${\mathscr{T}}\not\Rightarrow{\mathscr{W}}$ for arbitrary varieties (see [2, Example 4.4]). It is also known that ${\mathscr{W}}\Leftrightarrow{\mathscr{T}}$ for locally finite varieties (see [1, Theorem 9.6]). These facts, (i) ${\mathscr{W}}\Rightarrow{\mathscr{T}}$, (ii) ${\mathscr{T}}\not\Rightarrow{\mathscr{W}}$, (iii) ${\mathscr{W}}\Leftrightarrow{\mathscr{T}}$ for locally finite varieties, are the data supporting our statement that $\mathscr{W}$ is slightly (but properly) stronger than $\mathscr{T}$”.
@@
Now assume that Item (2) holds. Choose any ${\mathbf{A}}\in{\mathcal{V}}$ and any abelian tolerance $T$ of $\mathbf{A}$}. Our goal is to prove that $T$ is transitive, so choose elements satisfying aTbTca\;T\;b\;T\;c. Since $(a,b),(b,c)\in T$ and $T$ is a reflexive subalgebra of ${\mathbf{A}}^{2}$ we have $w^{T}((a,b),(b,b),(b,c))\in T$. Since $w(x,y,z)$ is a weak difference term for tolerances and $T$ is an abelian tolerance containing $(a,b)$ and $(b,c)$, we have $w^{T}((a,b),(b,b),(b,c))=(w^{{\mathbf{A}}}(a,b,b),w^{{\mathbf{A}}}(b,b,c))=(a,c)$, proving that $(a,c)\in T$. This shows that Item (3) holds.
[…]
For the congruences depicted in Figure 1, we have $\alpha,\alpha]=0$ and $\beta\leq\alpha$, so $\beta,\beta]=0$ by monotonicity of the commutator. In terms of the centralizer relation, this means ${\mathbf{C}}(\beta,\beta;0)$ holds. Now ${\mathbf{C}}(\beta,\beta;0)$ is equivalent to ${\mathbf{C}}(\beta,\beta;\delta)$ by Theorem 2.19 (4) above, so ${\mathbf{C}}(\beta,\beta;\delta)$ must also hold. We also have ${\mathbf{C}}(\delta,\beta;\delta)$ by Theorem 2.19 (8). The statements ${\mathbf{C}}(\beta,\beta;\delta)$ and ${\mathbf{C}}(\delta,\beta;\delta)$ may be combined to ${\mathbf{C}}(\beta\vee\delta,\beta;\delta)$ by Theorem 2.19 (5). By Theorem 2.19 (3) we derive that ${\mathbf{C}}(\beta\vee\delta,\delta\circ\beta\circ\delta;\delta)$ holds. If $S$ is the tolerance $\delta\circ\beta\circ\delta$ on $\mathbf{A}$}, then since $S\subseteq\beta\vee\delta$ we derive from ${\mathbf{C}}(\beta\vee\delta,\delta\circ\beta\circ\delta;\delta)$ and Theorem 2.19 (1) that ${\mathbf{C}}(S,S;\delta)$ holds. Now, by Theorem 2.19 (10), we derive that ${\mathbf{C}}(S/\delta,S/\delta;0)$ holds. But $T=S/\delta$, so ${\mathbf{C}}(T,T;0)$ holds. We have proved that $T$ is an abelian tolerance of ${\mathbf{A}}/\delta$.
@@
Theorem 3.24 of [3] is correct if the hypothesis that $\mathcal{V}$ has a Taylor term (or $\mathcal{V}$ satisfies a nontrivial idempotent Maltsev condition”) is strengthened to the hypothesis that $\mathcal{V}$ has a weak difference term.
@@ 4. Consequences
    ([3, Theorem 3.27]) This theorem examines the following property of a variety $\mathcal{V}$: given two perspective congruence intervals $\alpha/(\alpha\wedge\beta)$ and $(\alpha\vee\beta)/\beta$ of some ${\mathbf{A}}\in\mathcal{V}$, one interval is abelian if and only the other is abelian. Call this property $\mathscr{B}$. [3, Theorem 3.27] asserts that $\mathscr{T}$$\Rightarrow{\mathscr{B}}$. This implication is false. Using Corollary 3.2, we can correct Theorem 3.27 by strengthening the hypothesis $\mathscr{T}$ to the hypothesis $\mathscr{W}$. The corrected version then expresses that $\mathscr{W}$$\Rightarrow{\mathscr{B}}$. In fact, it is not hard to see that ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{B}})$. (See Theorem 5.1 (2) below.)
@@
    • All claims of all items in Theorem 4.16 are true when the hypothesis $\mathcal{V}$} has a Taylor term” is strengthened to $\mathcal{V}$} has a weak difference term”. (However, the two claims of 4.16 (3) were already known to be true in the presence of a weak difference term, cf. [5, Corollary 5.8].)
    We give more detail about the second bullet point. Using the labeling in Figure 1, it is stated in [3, Theorem 4.16 (2)] that a pentagon of congruences where ${\mathbf{C}}(\beta,\theta;\delta)$ holds cannot occur as a sublattice in a variety with a Taylor term. The proof given for this is correct. But then it is stated that a pentagon of congruences where ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ holds cannot occur as a sublattice. Call this property $\mathscr{C}$}. The proof given that ${\mathscr{T}}\Rightarrow{\mathscr{C}}$ depends on [3, Theorem 3.27], which (we have seen) is false without a weak difference term. In fact, it can be shown by arguments like those above that ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{C}})$. ((See Theorem 5.1 (3) below.)) To summarize, the part of [3, Theorem 4.16 (2)] that refers to ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ is false as stated, but becomes true when the hypothesis $\mathcal{V}$} has a Taylor term” is strengthened to $\mathcal{V}$} has a weak difference term”.
@@
[…]
@@
    ([3, Theorem 3.24]) is false as stated. Theorem 3.24 is true when the hypothesis $\mathcal{V}$} has a Taylor term” is strengthened to $\mathcal{V}$} has a weak difference term”.
  2. (2)
    ([3, Theorem 3.27]) is false as stated. Theorem 3.27 is true when the hypothesis $\mathcal{V}$} has a Taylor term” is strengthened to $\mathcal{V}$} has a weak difference term”.
@@
    • All claims of all items in Theorem 4.16 are true when the hypothesis $\mathcal{V}$} has a Taylor term” is strengthened to $\mathcal{V}$} has a weak difference term”.
@@
5. Afterword
We close this note by recording the results of the form ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{X}})$, where ${\mathscr{X}}={\mathscr{A}},{\mathscr{B}}$ or $\mathscr{C}$}, which we have considered in this paper, in the form of a theorem statement. Results of the form ${\mathscr{W}}\Leftrightarrow({\mathscr{T}}\;\&\;{\mathscr{X}})$, assert that Property $\mathscr{X}$ is Maltsev definable relative to $\mathscr{T}$ (= existence of a Taylor term) by the Maltsev condition that defines $\mathscr{W}$ (= existence of a weak difference term).
Observe that ${\mathscr{W}}~\Leftrightarrow~({\mathscr{T}}\;\&\;{\mathscr{X}})$ is equivalent to the conjunction of (i) ${\mathscr{W}}~\Rightarrow~{\mathscr{T}}$, (ii) ${\mathscr{W}}~\Rightarrow~{\mathscr{X}}$, and (iii) $({\mathscr{T}}\;\&\;{\mathscr{X}})~\Rightarrow~{\mathscr{W}}$. Implication (i) is known to be true: the class of varieties with a weak difference term is definable by a nontrivial idempotent Maltsev condition (cf. [4, Theorem 4.8]) and the class of varieties with a Taylor term is definable by the weakest nontrivial idempotent Maltsev condition (cf. Corollaries 5.2 and 5.3 of [6].) Implication (iii) may be rewritten in the form (iii)’ $({\mathscr{T}}\;\&\;\neg{\mathscr{W}})\Rightarrow\neg{\mathscr{X}}$. Altogether, we want to record why
  1. (ii)
    ${\mathscr{W}}~\Rightarrow~{\mathscr{X}}$, and
  2. (iii)’
    $({\mathscr{T}}\;\&\;\neg{\mathscr{W}})\Rightarrow\neg{\mathscr{X}}$
hold when ${\mathscr{X}}={\mathscr{A}},{\mathscr{B}}$ or $\mathscr{C}$}.
Theorem 5.1.
If $\mathcal{V}$ has a Taylor term, then any one of the following properties will hold for $\mathcal{V}$ if and only if $\mathcal{V}$ has a weak difference term.
  1. (1)
    ($\mathcal{V}$ satisfies $\mathscr{A}$}.) In any algebra of $\mathcal{V}$, the congruence generated by an abelian tolerance is abelian.
  2. (2)
    ($\mathcal{V}$ satisfies $\mathscr{B}$}.) Given two perspective congruence intervals $\alpha/(\alpha\wedge\beta)$ and $(\alpha\vee\beta)/\beta$ of some ${\mathbf{A}}\in\mathcal{V}$, one interval is abelian if and only the other is abelian.
  3. (3)
    ($\mathcal{V}$ satisfies $\mathscr{C}$}.) A pentagon of congruences, labeled as in Figure 2, cannot occur as a sublattice of $\operatorname{Con}({\mathbf{A}})$ if ${\mathbf{A}}\in\mathcal{V}$ and ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ holds.
    $\beta$$\delta$$\theta$$\alpha$$\beta\wedge\delta$
    Figure 2. Forbidden sublattice if ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ holds.
Proof.
Item (1) is Theorem 3.1 $(1)\Leftrightarrow(4)$.
It was noted above in Consequence (1) that the erroneous proof in [3] that $\mathscr{T}\Rightarrow\mathscr{B}$ is nevertheless valid as a proof of $\mathscr{W}\Rightarrow\mathscr{B}$. This gives us (ii) $\mathscr{W}$$\Rightarrow~{\mathscr{B}}$, as desired. To prove (iii)’, that is, to prove $({\mathscr{T}}\;\&\;\neg{\mathscr{W}})\Rightarrow\neg{\mathscr{B}}$, assume that $\mathcal{V}$ has a Taylor term but not a weak difference term. By Theorem 4.5 (5) of [2], there exists ${\mathbf{A}}\in\mathcal{V}$ with a pentagon in its congruence lattice labeled as in Figure 1. Since $\alpha,\alpha]=0$ in this pentagon, we get that $\beta,\beta]=0$ by monotonicity. Equivalently, the congruence quotient $\beta/(\beta\wedge\delta)$ is abelian. For the purpose of obtaining a contradiction assume that $\mathcal{V}$ satisfies $\mathscr{B}$. This allows us to deduce that the perspective interval $(\beta\vee\delta)/\delta=\alpha/\delta$ is also abelian. This is equivalent to the statement that ${\mathbf{C}}(\alpha,\alpha;\delta)$ holds, and therefore we derive that ${\mathbf{C}}(\beta,\theta;\delta)$ holds from Theorem 2.19 of [3] (i.e., by the monotonicity of the centralizer relation in the first two variables). But according to Theorem 4.2 of [2], no such pentagon can exist in the congruence lattice of any algebra with a Taylor term. This contradiction establishes Item (2).
To prove Item (3), we assert that that the erroneous proof in [3] of Theorem 4.16 (2) (i), which claims that $\mathscr{T}\Rightarrow\mathscr{C}$, is nevertheless valid as a proof of $\mathscr{W}\Rightarrow\mathscr{C}$. Indeed, the proof of $\mathscr{T}\Rightarrow\mathscr{C}$ from [3] is actually a valid proof that $\mathscr{B}\Rightarrow\mathscr{C}$ coupled together with an erroneous reference to $\mathscr{T}\Rightarrow\mathscr{B}$. We have argued above that $\mathscr{W}\Rightarrow\mathscr{B}$, so coupling this with the valid proof that $\mathscr{B}\Rightarrow\mathscr{C}$ yields $\mathscr{W}\Rightarrow\mathscr{C}$.
To complete the proof of this theorem, we must prove (iii)’ $({\mathscr{T}}\;\&\;\neg{\mathscr{W}})\Rightarrow\neg{\mathscr{C}}$. Assume that $\mathcal{V}$ fails $\mathscr{W}$, i.e. $\mathcal{V}$ does not have a weak difference term. By Theorem 4.5 (5) of [2] we know that there exists ${\mathbf{A}}\in\mathcal{V}$ with congruences labeled as in Figure 1 such that $\alpha,\alpha]=0=\beta\wedge\delta$ and ${\mathbf{C}}(\theta,\alpha;\delta)$. Since $\alpha,\alpha]=0=\beta\wedge\delta$ we have ${\mathbf{C}}(\alpha,\alpha;\beta\wedge\delta)$ by the definition of the commutator, hence ${\mathbf{C}}(\beta,\beta;\beta\wedge\delta)$ by the monotonicity of the centralizer in its first two variables. Thus, $\mathbf{A}$} has a pentagon of congruences that violate $\mathscr{C}$. This establishes that $\neg\mathscr{W}\Rightarrow\neg\mathscr{C}$, so the weaker implication $(\mathscr{T}\;\&\;\neg\mathscr{W})\Rightarrow\neg\mathscr{C}$ also holds. [^1] We have actually established that $\mathscr{W}\Leftrightarrow\mathscr{C}$.